A population of values has a normal distribution with μ=192.5μ=192.5 and σ=21.9σ=21.9. You intend to draw a random sample of size n=233n=233.

Find the probability that a single randomly selected value is between 190.1 and 194.4.
P(190.1 < X < 194.4) =

Find the probability that a sample of size n=233n=233 is randomly selected with a mean between 190.1 and 194.4.
P(190.1 < M < 194.4) =

Enter your answers as numbers accurate to 4 decimal places. Answers obtained using exact z-scores or z-scores rounded to 3 decimal places are accepted.
Answer #1

Solution-

According to question

A population of values has a normal distribution with

Mean = \mu =192.5 and

Standard deviation = \sigma =21.9

We have to draw a random sample of size n=233

Now,


(a)

The probability that a single randomly selected value is between 190.1 and 194.4 is

= P(190.1 <X< 194.4)

=P\left (\frac{190. 1-\mu}{\sigma}<z< \frac{194.4-\mu}{\sigma} \right )

=P\left ( \frac{190. 1-192.5}{21.9} <z<\frac{194.4-192.5}{21.9} \right )

P -2.4 21.9 <z< 1.9 21.9

=P\left ( -0.110<z<0.087 \right )

=P\left ( -0.110<z<0 \right )+P\left ( 0<z<0.087 \right )

= 0.0438 + 0.0347

= 0.0785

(b)

The probability that a sample of size n=233 is randomly selected with a mean between 190.1 and 194.4 is

= P(190.1 <Mean< 194.4)

=P\left ( \frac{190.1 -\mu}{\sigma/\sqrt{n}}<z<\frac{194.4-\mu}{\sigma/\sqrt{n}} \right )

=P\left ( \frac{190.1-192.5}{21.9/\sqrt{233}}<z<\frac{194.4-192.5}{21.9/\sqrt{233}} \right )

=P\left ( \frac{-2.4}{1.4347167}<z<\frac{1.9}{1.4347167} \right )

=P\left ( -1.673<z<1.324 \right )

=P\left ( -1.673<z<0 \right )+P\left ( 0<z<1.324 \right )

= 0.4528 + 0.40724

= 0.86004

0.8600 .